OpenStudy (anonymous):

A package of aluminum foil contains 50 ft^2 of foil, which weighs approximately 7.0 oz. Aluminum has a density of 2.70 g/cm^3. What is the approximate thickness of the foil in millimeters? Express your answer to two significant figures and include the appropriate units.

OpenStudy (anonymous):

So, how do I get from oz to g?

OpenStudy (anonymous):

Okay, so there are 198.4465 grams of aluminum.

OpenStudy (anonymous):

And 46451.52 square centimeters of aluminum.

OpenStudy (anonymous):

And the density is 2.70.

OpenStudy (anonymous):

mm=10cm. cm=V/A=(M/D)/A=M/DA=(198.4465)/((2.70)(46451.52))=0.0015822669248219154874523740817 That times 10 is 0.015822... But I only need 2 significant figures, so the answer is... 0.016 mm!